OpenStudy (anonymous):

What is the product of (2-i)(4+i)? And can you show me the steps of how to do it

OpenStudy (anonymous):

(2-i)(4+i) = 2(4+i) -i(4+i) = 8+2i-4i-i^2 = 8-2i-(-1) (because i^2=-1) = 8-2i+1 = 9-2i Hope that helped.

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

@bs07 You are welcomed.